HETAV PATEL

Grade 11,

A 0.02M solution of pyridinium hydrochloride has pH = 3.44. Calculate the ionization constant of pyridine ?

A 0.02M solution of pyridinium hydrochloride has pH = 3.44. Calculate the ionization constant of pyridine ?

Grade:12

1 Answers

Aryan Upadhyaya
6 years ago

pH = -log[H+]

[H+] = 3.63 x 10-4

Then, Kh =  3.63 x 10-4/0.02 = 6.63 x 10-5

Kh =  Kw/ Ka = 10-4/6.63 x 10-5 = 1.51 x 10-9